OpenStudy (anonymous):

find the next three numbers in the two patterns 1,4,16,64 3,6,12,24

OpenStudy (anonymous):

What do you think the pattern is for the first question?

OpenStudy (anonymous):

multiply last number by 4 in first sequence it is geomatric sequence as swcond one in second multiply 3 by 2 and then u get next number 6 then 6*2 =12 then 12*2=24 then 24*2=? then x *2 =??

OpenStudy (dlast):

1x4 =4 4x 4=16 16x4=64 64x4= ? ? x 4 = ? ? x 4 = ?